OpenStudy (anonymous):

Paul teaches the guitar in groups of three students. He rents a room for the guitar classes at $250 per month. He charges $30 per hour per student. Paul teaches six groups of students, and his net earnings per month are $1,910. Part A: If each student learns guitar for the same number of hours, write an equation that can be used to calculate the number of hours each student learns guitar in a month. (5 points) Part B: How many hours did each student learn guitar each month? Show your work and justify each step of your work. (5 points)

OpenStudy (anonymous):

@satellite73 Please help?

OpenStudy (anonymous):

@precal Help?

OpenStudy (anonymous):

@Ashleyisakitty please help?

OpenStudy (precal):

x is hours

OpenStudy (precal):

30x is what he is charging per student

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

he groups of three students tho

OpenStudy (anonymous):

teaches*

OpenStudy (precal):

6(3(30))x-250=1910

OpenStudy (anonymous):

ok so that answers part a right?

OpenStudy (anonymous):

@precal hello?

OpenStudy (precal):

my computer is acting weird

OpenStudy (anonymous):

oh ok hahah

OpenStudy (precal):

just solve the equation for part b, gotta go. My computer is acting really weird. It is hard for me to post anything.

OpenStudy (anonymous):

ok thanks tho

OpenStudy (anonymous):

@e.mccormick please help

OpenStudy (anonymous):

@amistre64 please can you help me?

OpenStudy (anonymous):

Paul teaches the guitar in groups of three students. Paul teaches six groups of students, How many students must there be, from these two sentences?

OpenStudy (anonymous):

hey @Chop-GTA5 sorry I'm here

OpenStudy (anonymous):

18 students

OpenStudy (anonymous):

@Chop-GTA5

OpenStudy (anonymous):

hold on sorry i have to figure this out again

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Yes. Now try writing the equation if x is the number of hours a student gets.

OpenStudy (anonymous):

18(30x) - 250 = 1910?

OpenStudy (anonymous):

18 students, 30 an hour per student, with x number of hours means 18*30*x

OpenStudy (anonymous):

540x - 250 = 1910

OpenStudy (anonymous):

yep now all you have to do is solve it

OpenStudy (anonymous):

when i solve it i will get the answer for part b right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then the equation is for part a

OpenStudy (anonymous):

ok let me solve it and tell u what I got so that i know i am not doing it wrong give me a sec

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

is it x = 4?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok great thank you very much! :D

OpenStudy (anonymous):

your very welcome :D

OpenStudy (anonymous):

Could you help me with just one more if u can?